OpenStudy (anonymous):

Solve the system. 2x+5y=33 x-y=6 (x, y) =

OpenStudy (anonymous):

y=3 x=9 multiply second equation by 2 and subtract simultaneously

OpenStudy (anonymous):

2x+5y=33 (-y )+ x-y=6 +(-y ) use additive inverse of y x=6+y | V 2(6+y) + 5y = 33 substitute 12+2y + 5y = 33 simplify (-12) + 12+7y = 33 + (-12) additive inverse of 12 1/7 [ 7y = 21 ] multiply by reciprocal of 7 y = 3 x= 6+ y use simplified equation x = 6 + (3 ) substitute x= 9 this is my best guess that this is correct